OpenStudy (anonymous):

divide and simplify 4√x^3/7√x

OpenStudy (anonymous):

Is this\[\frac {4}{\sqrt{x}} \div 7\sqrt{x}\]

OpenStudy (anonymous):

no 4√x^3 / 7√x

OpenStudy (anonymous):

here are my options A. 4√x^21 B. 4√x^2 C. √x^17 d. 28√x^17

OpenStudy (earthcitizen):

\[4*(x ^{3/2)-(1/2)})/7=(4 x)/7\]

OpenStudy (earthcitizen):

if those are your options, my guess is that the question is not clear enough can you re-write it with brackets ?

OpenStudy (anonymous):

ok (4√x^3) /( 7√x)

OpenStudy (earthcitizen):

then that is the answer

OpenStudy (anonymous):

\[\frac {4\sqrt{x^3}}{7\sqrt{x}} = \frac{4x}{7} \]

OpenStudy (angela210793):